OpenStudy (anonymous):

How many cubic millimeters (mm3) are in 2 cubic centimeters (cm3)? a. 1,000 mm3 b. 2,000 mm3 c. 20,000 mm3 d. 200 mm3

OpenStudy (anonymous):

\[2\text{ cm}^3=2\left(\text{ cm}\right)^3=2\left(10\text{ mm}\right)^3=2\left(1000\text{ mm}^3\right)=2000\text{ mm}^3\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

yw